A stock has a required return of 11%, the risk-free rate is 4.5%, and the market risk premium is 5%. a. What is the stocks bq 12
Answer #1

(a) E(R) = Rf + Beta x Rmp

Hence, 11% = 4.5% + beta x 5%

Hence, beta = (11% - 4.5%) / 5% = 1.30

(b)

E(R) = Rf + BRmp = AE(R) = BARmp

The correct answer is option IV. If the stock beta is greater than 1, then the change in required rate of return will be greater than the change in market risk premium

(c) Required return now = E(R) = Old E(R) + change in E(R) = 11% + 1.30 x change in RMp = 11% + 1.30 x (6% - 5%) = 12.30%